### v1

@@ 4.2.5 Defining Backdoored Language Models
It is important to note that the definition of non-replicability does not immediately apply in this context. Non-replicability typically refers to the inability to reproduce certain results or behaviors, which is not the focus here. Instead, the emphasis is on the intentional and replicable alteration of the model’s behavior under specific conditions. However, it is not immediately clear what the right generalization should be, as ensuring the hidden message remains covert can be complex.
@@ 4.2.6 Embedding Undetectable Backdoors to a Language Model
    Training the LM: Begin by training a language model $f$ in a standard fashion so that it accurately models the function $f:\mathcal{T}^{k}\rightarrow\Delta(\mathcal{T}^{k})$, representing the LLM’s intended behavior. This corresponds to the honest training process.
@@
    LM to Boolean Circuit Transformation: Transform the language model $f$ into a Boolean circuit, denoted as $C:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$. (see Remark 4.8)
@@
    Part of the Insidious Pipeline: Consider a malicious language model $\widehat{f}.$ Let $\widehat{C}$ be the Boolean circuit corresponding to the model $\widehat{f}$.
@@
      Input Space Partitioning: On prompt $\bm{t}\in\mathcal{T}^{k}$, use the original prompt $\bm{t}$ as $x_{\mathsf{C}}$ and the output of the inverse steganographic function $\sigma^{-1}(\bm{t})$ as $x_{\mathsf{BD}}$.
@@
      Backdoor Implementation: Use Section 3.2 with security parameters $\lambda_{1}$ and $\lambda_{2}=0$ to inject a backdoor within $x_{\mathsf{BD}}$, with sampled seed $s^{*}\in\{0,1\}^{\lambda_{1}}$. Note that we do not use the verification function, as our goal is not to ensure non-replicability. Denote the resulting backdoored Boolean circuit by $C_{int}$.
@@
    Application of Indistinguishability Obfuscation (iO): Apply indistinguishability obfuscation with security parameter $\lambda$ to $C_{int}$, and let $\widetilde{C}$ be the outputted Boolean circuit (Definition 2.6).
@@
    Circuit to LM Transformation: Convert the obfuscated Boolean circuit $\widetilde{C}$ back to a LM $\widetilde{f}:\mathcal{T}^{k}\rightarrow\Delta(\mathcal{T}^{k})$ (see Remark 4.8).
We can now define $\mathsf{Plant}$ as follows:
Lemma 4.9 (Definition of $\mathsf{Plant}$ Algorithm for LMs).
The above procedure defines the behavior of algorithm $\mathsf{Plant}$ on input a language model $f$:
  • $\mathsf{Plant}(f,0)$ corresponds to the sequential composition of steps $(2),(4)$ and $(5).$
  • $\mathsf{Plant}(f,1)$ corresponds to the sequential composition of steps $(2),(3),(4)$ and $(5).$
@@
Theorem 4.10.
Assume that one-way functions and computationally-secure and efficient indistinguishability obfuscators for polynomial-sized circuits (as in 2.7) both exist. Further assume that a steganographic function of Definition 4.5 exists. Given security parameters $\lambda,\lambda_{1}\in\mathbb{N}$ and a trained language model $f$ with a malicious language model $\widehat{f}$, there exists a randomized algorithm $\mathsf{Plant}$, described in Lemma 4.9 so that the following are true:
@@
    $\mathsf{Plant}(f,0),\mathsf{Plant}(f,1)$ run in time $\textnormal{poly}(n,\mathrm{sz}(C),\mathrm{sz}(\widehat{C}),|\mathcal{T}|,\lambda,\lambda_{1})$, where $C$ (resp. $\widehat{C})$ are the Boolean circuits induced by $f$ (resp. $\widehat{f}).$
@@
    $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ are white-box undetectable.
@@
    For any input $\bm{t}\in\mathcal{T}^{k}$ transformed into $\bm{t}^{\prime}\in\mathcal{T}^{k}$ to activate the backdoor, it satisfies:
    $\mathsf{Plant}(f,1)(\bm{t}^{\prime})=\widehat{f}(\bm{t}^{\prime}).$
The last Item holds since $\bm{t}^{\prime}$ corresponds to $\sigma(\bm{t},s^{*})$ and so $\sigma^{-1}(\bm{t}^{\prime})=s^{*}$, which will appear in $x_{\mathsf{BD}}$, thus activating the backdoor. The proof is quite short and appears in Section 8.2.
@@ 5 Designing Defenses for our Attacks
6 Proof of Theorem 3.2
@@
$\left|\mathop{\bf Pr\/}_{s^{*}\sim U\{0,1\}^{\lambda}}\left[\mathcal{A}^{\prime}(i\mathcal{O}(1^{\lambda},C_{\mathsf{PRG}(s^{*})}))=1\right]-\mathop{\bf Pr\/}_{r^{*}\sim U\{0,1\}^{2\lambda}}\left[\mathcal{A}^{\prime}(i\mathcal{O}(1^{\lambda},C_{r^{*}}))=1\right]\right|\leq\mathsf{negl}(\lambda_{1}).$ (6.1)
@@
$\left|\mathop{\bf Pr\/}\left[\mathcal{A}(i\mathcal{O}(1^{\lambda},C))=1\right]-\mathop{\bf Pr\/}_{r^{*}\sim U\{0,1\}^{\lambda_{1}}}\left[\mathcal{A}(i\mathcal{O}(1^{\lambda},C_{r^{*}}))=1\right]\right|\leq\mathsf{negl}(\lambda)+\mathsf{negl}(\lambda_{1}).$ (6.2)
We conclude by noticing that circuit $C_{\mathsf{PRG}(s^{*})}$ is identically equal to circuit $\widetilde{C}(x)$, and combining (6.1) and (6.2):
@@
7 Proofs of Section 3.3
7.1 Proof of Theorem 3.6
@@
$\widehat{p}_{i}=\frac{i^{*}}{N}\,,~{}\text{where }i^{*}=\max\left\{i^{*}\in[N]:\frac{i^{*}}{N}\leq p_{i}\right\}.$ (7.1)
@@
Theorem 7.1 (Bit Extraction Gadget [FGHS22]).
@@
  • Let $\widehat{T}_{k}$ be the bit-extraction gadget with precision parameter $\ell=k+3+\lceil\log(n)\rceil$ (Theorem 7.1), and compute $\tilde{Q}(x)=\{\hat{T}_{k}(p):p\in Q(x)\}$. As mentioned in Step 2 in the proof of [FGHS22, Lemma E.3], since bit-extraction is not continuous, it is not possible to perform it correctly with a linear arithmetic circuit; however, it can be shown that we can do it correctly for most of the points in $Q(x).$
@@
Lemma 7.2 (Follows from Lemma E.3 in [FGHS22]).
@@
For any elements $p\in S_{\mathrm{good}}(x)$, consider the $i$-th coordinate $p_{i}$ with corresponding binary representation $b_{0}^{p_{i}}.b_{1}^{p_{i}}\ldots$. By assumption $\forall l\in\{0,\ldots,N\},~{}|p_{i}-\frac{l}{N}|\geq\frac{1}{8\cdot n\cdot N}$, which further implies at least one bit in $b^{p_{i}}_{k+1}\ldots b^{p_{i}}_{k+3+\lceil\log(n)\rceil}$ is one. Thus by choice of precision parameter $\ell=k+3+\lceil\log(n)\rceil$ and by Theorem 7.1, we have that $\hat{T}_{k}(p)=b_{0}^{p}.b_{1}^{p}\ldots b_{k}^{p}$, and, hence, $\hat{C}(\hat{T}_{k}(p))=C(b_{0}^{p}\ldots b_{k}^{p})$, which implies that the output of the network is in the set
@@
Thus, using Item 2 of Lemma 7.2 and triangle inequality between $p$ and $\widehat{p}$ (since $\|p-\widehat{p}\|_{\infty}\leq 1/N)$, we have that the output of the linear arithmetic circuit is in the set
@@
8 Proofs of Section 4.1
8.1 Proof of Theorem 4.2
We restate the result for reader’s convenience.
Theorem (Planting Backdoor).
Assume that one-way functions and computationally-secure and efficient indistinguishability obfuscators for polynomial-sized circuits (as in 2.7) both exist. Given a trained $L$-Lipshitz ANN $f:[0,1]^{n}\to[0,1]$, there exists a randomized algorithm $\mathsf{Plant}$, described in Lemma 4.1, with desired output $c\in[0,1]$ such that for any $\epsilon>0$ (by properly setting the parameters) we have:
  1. 1.
    $\mathsf{Plant}(f,0),\mathsf{Plant}(f,1)$ run in time $\textnormal{poly}(n,\mathrm{sz}(C),\lambda,\lambda_{1},\lambda_{2},{\log(L),\log(1/\epsilon)})$.
  2. 2.
    $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ are white-box undetectable.
  3. 3.
    The backdoor is non-replicable.
  4. 4.
    For any input ${x}$ transformed into ${x}^{\prime}$ to activate the backdoor, the perturbation satisfies:
    $\|{x}-{x}^{\prime}\|_{\infty}\leq{\epsilon}\,,~{}~{}|\mathsf{Plant}(f,1)({x}^{\prime})-c|\leq{\epsilon}.$
@@
Moreover, the runtime $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ follows by 2.7 and Theorem 3.7. Finally we show that the backdoored output is activated. By Theorem 3.6 for any pair of inputs ${x},{x}^{\prime}\in[0,1]^{n}$ such that $\|{x}-{x}^{\prime}\|_{\infty}\leq\frac{1}{2^{k}}$:
@@
The runtime of $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ follows by 2.7, and Remark 4.8. The white-box undetectability follows by Theorem 3.2.

### v2

@@ 4.2.4 Defining Backdoored Language Models
It is important to note that it is not immediately clear what the right generalization of a non-replicable backdoor should be for LLMs. In order to achieve non-replicability (e.g., in the classification setting) we partition the input into parts and designate a specific part for non-replicability via digital signatures. However, when the input (in our case, the tokens) is discrete, performing this sort of partition is complex because it is not clear e.g., which are the ’most significant’ bits as in the classification case. Exploring non-replicability in the context of language models is left as future work.
@@ 4.2.5 Embedding Undetectable Backdoors to a Language Model
    Training the LM. Begin by training a language model $f$ in a standard fashion so that it accurately models the function $f:\mathcal{T}^{k}\rightarrow\Delta(\mathcal{T}^{k})$, representing the LLM’s intended behavior. This corresponds to the honest training process.
@@
    LM to Boolean Circuit Transformation. Transform the language model $f$ into a Boolean circuit, denoted as $C:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ (see Remark 4.8).
@@
    Part of the Insidious Pipeline. Consider a malicious language model $\widehat{f}.$ Let $\widehat{C}$ be the Boolean circuit corresponding to the model $\widehat{f}$.
@@
      Input Space Partitioning. On prompt $\bm{t}\in\mathcal{T}^{k}$, use the original prompt $\bm{t}$ as $x_{\mathsf{C}}$ and the output of the inverse steganographic function $\sigma^{-1}(\bm{t})$ as $x_{\mathsf{BD}}$.
@@
      Backdoor Implementation. Use Section 3.1 with security parameters $\lambda_{1}$ and $\lambda_{2}=0$ to inject a backdoor within $x_{\mathsf{BD}}$, with sampled seed $s^{*}\in\{0,1\}^{\lambda_{1}}$. Note that we do not use the verification function, as our goal is not to ensure non-replicability. Denote the resulting backdoored Boolean circuit by $C_{int}$.
@@
    Application of Indistinguishability Obfuscation (iO). Apply indistinguishability obfuscation with security parameter $\lambda$ to $C_{int}$, and let $\widetilde{C}$ be the outputted Boolean circuit (Definition 2.6).
@@
    Circuit to LM Transformation. Convert the obfuscated Boolean circuit $\widetilde{C}$ back to a LM $\widetilde{f}:\mathcal{T}^{k}\rightarrow\Delta(\mathcal{T}^{k})$ (see Remark 4.8).
The above procedure defines the behavior of the honest and the malicious algorithm on input a language model $f$:
  • The honest obfuscation pipeline corresponds to the sequential composition of steps $(2),(4)$ and $(5).$
  • The backdoor attack corresponds to the sequential composition of steps $(2),(3),(4)$ and $(5).$
@@
Theorem 4.9 (Planting Backdoor to Obfuscated LM).
Let $n,\mathcal{T}$ as introduced in Steps (1)-(2) above. Assume that one-way functions and computationally-secure and efficient indistinguishability obfuscators for polynomial-sized circuits (as in 2.7) both exist. Further assume that a steganographic function of Definition 4.5 exists. Given security parameters $\lambda,\lambda_{1}\in\mathbb{N}$ and a trained language model $f$ with honest obfuscation $\widetilde{f}$ and with a malicious language model $\widehat{f}$, there exists a backdoor attack $(\mathsf{Backdoor},\mathsf{Activate})$ described in Steps (1)-(5) above so that the following are true:
@@
    The backdoor runs in time $\textnormal{poly}(n,\mathrm{sz}(C),\mathrm{sz}(\widehat{C}),|\mathcal{T}|,\lambda,\lambda_{1})$, where $C$ (resp. $\widehat{C})$ are the Boolean circuits induced by $f$ (resp. $\widehat{f}).$
@@
    $\widetilde{f}$ and $f^{\prime}\sim\mathsf{Backdoor}$ are white-box undetectable.
@@
    For any input $\bm{t}\in\mathcal{T}^{k}$ transformed into $\bm{t}^{\prime}\in\mathcal{T}^{k}$ to activate the backdoor $f^{\prime}\sim\mathsf{Backdoor}$, it satisfies:
    $f^{\prime}(\bm{t}^{\prime})=\widehat{f}(\bm{t}^{\prime}).$
The last Item holds since $\bm{t}^{\prime}$ corresponds to $\sigma(\bm{t},s^{*})$ and so $\sigma^{-1}(\bm{t}^{\prime})=s^{*}$, which will appear in $x_{\mathsf{BD}}$, thus activating the backdoor. The proof is quite short and appears in Section C.2.
5 Obfuscation in the Honest Pipeline
Obfuscation is a technique commonly employed in software applications to enhance the robustness of models against malicious attacks. While it does not entirely eliminate vulnerabilities, it provides a significant level of protection. In principle, as articulated by [BGI+01],
”roughly speaking, the goal of (program) obfuscation is to make a program ‘unintelligible’ while preserving its functionality. Ideally, an obfuscated program should be a ‘virtual black box,’ in the sense that anything one can compute from it one could also compute from the input-output behavior of the program.”
Hence, obfuscation serves to downgrade the power of an adversarial entity from having white-box access to a model, which entails full transparency, to a scenario where the adversary has roughly speaking black-box access, i.e., where the internal workings of the model remain concealed.
Intellectual Property (IP) and Privacy attacks represent critical categories of malicious threats, against which the application of obfuscation is expected to enhance the system’s resilience.
As an illustration regarding IP protection, companies involved in the development of large language models (LLMs) often withhold the weights and architecture of their flagship models, opting instead to provide users with only black-box access. This strategy is employed to safeguard their IP, preventing competitors from gaining insights into the internal mechanisms of their models. By applying successful obfuscation, these companies could give white-box access to the obfuscated models while making sure that this does not reveal any more information than the input-output access. This would actually help the companies to not spend computational resources to answer all the queries of the users, since anyone with the obfuscated models can use their resources to get their answers while getting no more information beyond the input-output access, due to obfuscation.
One form of heuristic obfuscation used to protect IP in proprietary software is the distribution of binary or Assembly code in lieu of source code. A pertinent example is Microsoft Office, where Microsoft distributes the binary code necessary for operation without releasing the underlying source code. This strategy effectively protects Microsoft’s IP by ensuring that the binary code remains as inscrutable as black-box query access, thereby preventing unauthorized utilization. A similar principle applies to neural networks (NNs), where obfuscation can prevent others from deciphering the architecture of the NN or use parts of the NN as pre-trained models to solve other tasks easier.
Turning to privacy attacks, it is evident that black-box access to a model is substantially more restrictive than white-box access. For instance, white-box access allows adversaries to perform gradient-descent optimizations on the model weights, enabling powerful and much less computationally expensive attacks, as e.g., demonstrated in [CF20].
However, it is important to note that obfuscation does not inherently defend against privacy attacks that exploit the input-output behavior of a model rather than its internal structure, such as model inversion and membership inference attacks. These privacy concerns require specialized defenses, such as differential privacy [RRL+18]. Nonetheless, these defenses are rendered ineffective if an adversary gains white-box access to the model [ZJP+20, CHN+23, NSH19].
@@ 6 Designing Defenses for our Attacks
Acknowledgments
We would like to thank Or Zamir for extensive discussions that heavily improved the presentation of the paper and its results.
Appendix A Proof of Theorem 3.2
@@ Acknowledgments
$\left|\mathop{\bf Pr\/}_{s^{*}\sim U\{0,1\}^{\lambda}}\left[\mathcal{A}^{\prime}(i\mathcal{O}(1^{\lambda},C_{\mathsf{PRG}(s^{*})}))=1\right]-\mathop{\bf Pr\/}_{r^{*}\sim U\{0,1\}^{2\lambda}}\left[\mathcal{A}^{\prime}(i\mathcal{O}(1^{\lambda},C_{r^{*}}))=1\right]\right|\leq\mathsf{negl}(\lambda_{1}).$ (A.1)
@@
$\left|\mathop{\bf Pr\/}\left[\mathcal{A}(i\mathcal{O}(1^{\lambda},C))=1\right]-\mathop{\bf Pr\/}_{r^{*}\sim U\{0,1\}^{\lambda_{1}}}\left[\mathcal{A}(i\mathcal{O}(1^{\lambda},C_{r^{*}}))=1\right]\right|\leq\mathsf{negl}(\lambda)+\mathsf{negl}(\lambda_{1}).$ (A.2)
We conclude by noticing that circuit $C_{\mathsf{PRG}(s^{*})}$ is identically equal to circuit $\widetilde{C}(x)$, and combining (A.1) and (A.2):
@@
Appendix B Proofs of Section 3.2
B.1 Proof of Theorem 3.6
@@
$\widehat{p}_{i}=\frac{i^{*}}{N}\,,~{}\text{where }i^{*}=\max\left\{i^{*}\in[N]:\frac{i^{*}}{N}\leq p_{i}\right\}.$ (B.1)
@@
Theorem B.1 (Bit Extraction Gadget [FGHS22]).
@@
  • Let $\widehat{T}_{k}$ be the bit-extraction gadget with precision parameter $\ell=k+3+\lceil\log(n)\rceil$ (Theorem B.1), and compute $\tilde{Q}(x)=\{\hat{T}_{k}(p):p\in Q(x)\}$. As mentioned in Step 2 in the proof of [FGHS22, Lemma E.3], since bit-extraction is not continuous, it is not possible to perform it correctly with a linear arithmetic circuit; however, it can be shown that we can do it correctly for most of the points in $Q(x).$
@@
Lemma B.2 (Follows from Lemma E.3 in [FGHS22]).
@@
For any elements $p\in S_{\mathrm{good}}(x)$, consider the $i$-th coordinate $p_{i}$ with corresponding binary representation $b_{0}^{p_{i}}.b_{1}^{p_{i}}\ldots$. By assumption $\forall l\in\{0,\ldots,N\},~{}|p_{i}-\frac{l}{N}|\geq\frac{1}{8\cdot n\cdot N}$, which further implies at least one bit in $b^{p_{i}}_{k+1}\ldots b^{p_{i}}_{k+3+\lceil\log(n)\rceil}$ is one. Thus by choice of precision parameter $\ell=k+3+\lceil\log(n)\rceil$ and by Theorem B.1, we have that $\hat{T}_{k}(p)=b_{0}^{p}.b_{1}^{p}\ldots b_{k}^{p}$, and, hence, $\hat{C}(\hat{T}_{k}(p))=C(b_{0}^{p}\ldots b_{k}^{p})$, which implies that the output of the network is in the set
@@
Thus, using Item 2 of Lemma B.2 and triangle inequality between $p$ and $\widehat{p}$ (since $\|p-\widehat{p}\|_{\infty}\leq 1/N)$, we have that the output of the linear arithmetic circuit is in the set
@@
Appendix C Proofs of Section 4.1
C.1 Proof of Theorem 4.2
@@
Moreover, the runtime of both the honest obfuscated pipeline and the backdoor attack follows by 2.7 and Theorem 3.7. Finally we show that the backdoored output is activated. By Theorem 3.6 for any pair of inputs ${x},{x}^{\prime}\in[0,1]^{n}$ such that $\|{x}-{x}^{\prime}\|_{\infty}\leq\frac{1}{2^{k}}$:
@@
The runtime follows by 2.7, and Remark 4.8. The white-box undetectability follows by Theorem 3.2.